# Checklists Are Better Than Reward Models For Aligning Language Models

### v1

@@ Abstract
Language models must be adapted to understand and follow user instructions. Reinforcement learning is widely used to facilitate this – typically using fixed criteria such as “helpfulness” and “harmfulness”. In our work, we instead propose using flexible, instruction-specific criteria as a means of broadening the impact that reinforcement learning can have in eliciting instruction following. We propose “Reinforcement Learning from Checklist Feedback” (RLCF). From instructions, we extract checklists and evaluate how well responses satisfy each item—using both AI judges and specialized verifier programs—then combine these scores to compute rewards for RL. We compare RLCF with other alignment methods applied to a strong instruction following model (Qwen2.5-7B-Instruct) on five widely-studied benchmarks – RLCF is the only method to improve performance on every benchmark, including a 4-point boost in hard satisfaction rate on FollowBench, a 6-point increase on InFoBench, and a 3-point rise in win rate on Arena-Hard. These results establish checklist feedback as a key tool for improving language models’ support of queries that express a multitude of needs. [^1] We plan on releasing our models, our dataset of checklists (WildChecklists), and code to the public shortly.
@@ 1 Introduction
Language models must follow user instructions to be useful. As the general public integrates language model-based assistants into their completion of daily tasks, there is an expectation that language models can faithfully follow the users’ requests (Liu et al., 2024a). As users develop more confidence in models’ ability to fulfill complex requests, these models are increasingly given rich, multi-step instructions that require careful attention to specifications (Zhao et al., 2024; Zheng et al., ).
@@
Today’s models are almost universally trained to follow instructions via a two-step process: instruction finetuning, followed by reinforcement learning from human feedback (RLHF). Instruction finetuning, where the model is trained to mimic responses generated by annotators (Raffel et al., 2019), has historically been the primary workhorse for imbuing language models with some amount of instruction-following ability (Wang et al., 2022; Chung et al., 2022; Xu et al., 2024; Lambert et al., 2024a). Model developers then frequently employ RLHF, where the model is trained to generate responses that look more like labeled “good” responses than “bad” responses, as a refinement step to decrease the likelihood that the model exhibits predefined poor behaviors (typically harmful behaviors) (Ziegler et al., 2019; Bai et al., 2022). Unlike “verifiable” tasks where reinforcement learning is a workhorse (DeepSeek-AI et al., 2025; Lambert et al., 2024a; Pyatkin et al., 2025), reinforcement learning remains difficult to utilize for ambiguous or “non-verifiable” tasks, such as instruction following. What would it take to make RL a general-purpose solution at eliciting desirable behaviors in subjective or open-ended settings?
@@
Figure 1: RL on Checklist Feedback consistently improves Qwen2.5 7B Instruct, whereas every other source of automatic feedback gives mixed results.
@@
We believe the solution must involve producing better reward signals. Recent work on RL for language model alignment has focused on automatically obtaining feedback on model behavior, either by (1) exclusively using instructions with verifiable answers (Dong et al., 2024; Pyatkin et al., 2025), (2) grading responses with specially-trained reward models (Wang et al., 2024a; Eisenstein et al., 2023), or (3) distilling preferences from a larger prompted model (Bai et al., 2022; Tunstall et al., 2023). Using instructions with verifiable answers restricts the aspects of response quality that can be learned to exact answer correctness or syntax/format adherence (ignoring other qualities, e.g. topicality or style). Specially-trained reward models are powerful, but their notion of rewards can be arbitrary, leading to reward hacking (Eisenstein et al., 2023). When distilling preferences from a larger prompted LM, that LM is challenged with deciding what aspects to consider when grading a response, reducing the so-called “generator-verifier gap” that enables RL (Swamy et al., 2025). Even if multiple prompts are written to capture values of interest, this assumes that a fixed set of criteria can be comprehensive (Bai et al., 2022; Glaese et al., 2022a).
@@
In this paper, we ask: “how can we grade responses to instructions in a manner that is automatic (requires no human annotation), flexible (considers all aspects of response quality), intuitive (aligned with perceptible differences in responses), and applicable to any instruction or response, to enable more effective use of RL in language model alignment?” As an answer, we propose extracting dynamic checklists from instructions – an approach we term Reinforcement Learning from Checklist Feedback (RLCF). This approach makes it likely that our evaluation focuses on flexible lists of distinct criteria while also reducing the problem of grading responses to answering a series of specific yes/no questions, which can be answered by an AI judge or by executing a verification program.
@@
On 5 benchmarks covering both constrained instruction-following (IFEval, InFoBench, FollowBench) and general conversational assistance (AlpacaEval, Arena-Hard), we find that RLCF provides benefits on all instruction following benchmarks while maintaining improved performance on general conversational assistance benchmarks. In contrast, all alternative forms of AI feedback lead to mixed results, as shown in Figure 1. RLCF provides a 5.4% relative improvement over Qwen2.5-7B-Instruct in average hard satisfaction rate on FollowBench, a 6.9% relative improvement in overall requirement following ratio on InFoBench, and a 6.4% relative improvement on Arena-Hard (Jiang et al., 2023; Qin et al., 2024; Li et al., 2024). Despite these considerable improvements, RLCF simply requires a teacher model, with no need for additional data or human annotations, making this approach amenable to diverse languages or domains. We provide evidence that checklist-based rewards are well-correlated to human preference judgments (comparable to many finetuned reward models) while providing a stronger learning signal than alternatives.
@@ 2 Checklist Generation
To satisfy definition #3, checklists must be comprehensive (cover most relevant aspects of quality) and natural (entailed by their corresponding instructions). Based on the observation that false positive rewards are often more detrimental to reinforcement learning than false negatives (Huang et al., 2024), we want checklists that are objective (facilitate automatic verification) and atomic (each requirement focuses on a single aspect of quality), to make requirement checking easier.
@@
  • Direct: We simply prompt an LM to extract a checklist from a given instruction (Cook et al., 2024). This approach is intuitive and simple but risks repeating the original instruction via these individual criteria, which may limit comprehensiveness and objectiveness.
@@
To compare these, we generate checklists for all instructions in InFoBench (Qin et al., 2024). We use gpt-4o to blindly evaluate each of these checklists on naturalness, objectivity, comprehensiveness, and atomicity, then select the better one overall. We manually perform the same evaluation on a subset of 50 instructions from the “Easy Set” of InFoBench.
@@
Regularization via universal criteria. In initial experiments, we found that optimizing for checklist completion led models to sometimes generate high-level overviews of the response instead of the intended response, suggesting reward hacking. In prior work, Sun et al. (2023) reported a similar issue when training models with instructable reward models, addressed by adding three manually-chosen instructions to their reward model in all cases. Following this and other works that perform RL using global principles (Glaese et al., 2022b; Bai et al., 2022), we added one “universal requirement” to all generated checklists. This universal requirement stated “Does the response satisfy the following two criteria: 1) The response directly address the request without excessive or off-topic information not necessary for addressing the user’s instruction? 2) The response should match the context and the instruction, whether it requires professionalism, friendliness, formality, or neutrality.”, with a corresponding importance weight of 100/100.
@@
Dataset Generation Using the candidate-based method, we generate checklists for 130,000 instructions from WildChat to create a new dataset, WildChecklists. To generate candidate responses for our method, we use Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B, and Qwen2.5-7B (Yang et al., 2024). Qwen2.5-72B-Instruct is the checklist generator model for both methods.
@@ 3 Reinforcement Learning from Checklist Feedback
Sampling Candidate Responses. To facilitate offline RL, we first sample response pairs from our base policy. For each prompt, we sample two responses with a temperature of 1.3 and a top-p of 0.9 (Holtzman et al., 2019). This is simpler than prior works on RL-based language model alignment that synthetically perturb prompts to induce greater complexity (Sun et al., 2024; Dong et al., 2024).
Flexible Scoring Given a prompt, a response, and an individual checklist item, we use a combination of an LM judge and a verifier program to grade the response. For each checklist item, the judge model (Qwen2.5-72B-Instruct) generates a numerical score between 0 and 100. The prompt we use for grading responses is shown in Appendix B. To reduce the variance of this score, we sample 25 numerical scores from the model, and we then take the average of these 25 scores [^2] We sample responses using the n parameter in vLLM (Kwon et al., 2023). This approach follows prior work that describes the importance of using the mean score rather than mode score from an LM-as-a-judge model (Wang et al., 2025). Regardless, this makes the AI judge component the computational bottleneck of our pipeline. In Section 5.5, we show that n can be significantly reduced, at a modest accuracy cost. .
When applicable, we also use a verifier program to perform grading. LLMs struggle to evaluate criteria that measure discrete properties of text, such as “does the response contain the letter R at least three times?” or “does the response contain one of the following keywords […]?” (Fu et al., 2024). To better handle such constraints, we follow prior work in generating a verification program when applicable (Dong et al., 2024; Zhou et al., 2023). Our prompt, listed in Appendix B, asks the model to produce code only when the model is highly confident the requirement can be exactly verified with a program. If the program successfully processes a response string, we convert the Boolean result to an integer (0 or 100), which is averaged with the AI judge’s score. [^3] This approach is much simpler than the most relevant prior work that uses programs to evaluate responses, AutoIF (Dong et al., 2024), which uses test-case generation and LM-based filters to remove low-quality programs.
Preference Tuning. Given a separate numerical score for each criterion for each response, we take the weighted average of these scores, weighting by the importance score generated for each criterion. To produce a more informative learning signal, we keep only the 40% of response pairs with the greatest difference along at least one criterion of its corresponding checklist. This removes response pairs that are too similar in quality to offer a useful reward signal. We then assign the higher scoring response as “chosen” and the lower as “rejected”, and we use these as a preference pair for direct preference optimization (Rafailov et al., 2023).
@@ 4.1 Experimental Details
Training Data As a fixed source of instructions for all methods, we use WildChat, a set of natural conversations between users and AI language models crowdsourced from users across the world (Zhao et al., 2024). We filter out conversations that are non-English, toxic, or longer than two turns.
@@
Training We finetune the model for 2 epochs using DPO with a batch size of 1024 and a maximum sequence length of 2048. We use a cosine learning rate schedule with a max LR of 3e-6 and a min LR of 2e-6444When training models with Ultrafeedback, we instead used a minimum learning rate of 3e-7. We found this parameter resulted in a slightly stronger baseline when learning from this feedback.. We use OpenRLHF for training (Hu et al., 2024), and we train on one 8xH100 node with 80GB GPU memory, which took roughly 3 hours for each model.
Benchmark Data We evaluate our method on five benchmarks: IFEval (Zhou et al., 2023), InFoBench (Qin et al., 2024), FollowBench (Jiang et al., 2023), AlpacaEval (Dubois et al., 2024), and Arena-Hard (Li et al., 2024). The first three of these measure instruction following ability in the presence of fine-grained constraints. The last two measure “general-purpose” instruction-following ability, using naturalistic instructions based on user queries collected in the wild.
@@ 4.2 Baselines
Instruction Finetuning: We compare with instruction finetuning, to isolate the benefit of additional knowledge from the manner it is given (ground truth or rewards). Here, we distill (Hinton et al., 2015) from a larger model, Qwen2.5-72B-Instruct, finetuned via LlamaFactory (Zheng et al., 2024).
Reward Models: We mirror our training approach for learning from checklist feedback, but using state-of-the-art reward models to decide which response should be chosen or rejected. Here, we keep the 40% of prompts and responses with the greatest difference in scalar rewards. We consider following reward models as baselines (Skywork/Skywork-Reward-Gemma-2-27B) (Liu et al., 2024b) and ArmoRM-Llama3-8B-v0.1) (Wang et al., 2024b). Both are highly rated on RewardBench (Lambert et al., 2024b)555 Skywork/Skywork-Reward-Gemma-2-27B and ArmoRM-Llama3-8B-v0.1 are ranked as #4 and #24, respectively, on RewardBench as of July 2025., and ArmoRM has been very effective for alignment in prior work (Meng et al., 2024).
Prompted AI Judge: We lastly compare against using the same “teacher” model as a judge, without using checklists. We query this teacher in two settings: 1) “Ultrafeedback”, where the judge rates all candidate responses from 1-5 (Cui et al., 2023) separately across four quality aspects (instruction following, helpfulness, truthfulness, honesty) and averages these scores; and 2) AI Judge, where a near-identical prompt as RLCF is used (§3) to similarly sample 25 scores between 0 and 100 from the judge. This uses the AI judge the same way as RLCF, just without a checklist.
@@ 5.1 RL from Checklist Feedback consistently improves language models
Table 2: RLCF improves performance modestly on a format-based constrained instruction following benchmark (IFEval) and significantly on an open-ended constrained instruction following benchmark (InFoBench). RL on rewards from off-the-shelf reward models help on InFoBench but hurt on IFEval. We show positive results (relative to the baseline) in blue, negative in orange, and neutral (within 0.5) in gray; the top variant of a given model is bolded.
@@
Table 3: RLCF leads to significant improvements on FollowBench uniformly across all metrics when starting with an instruction-tuned model, while using an off-the-shelf reward model for preference labeling leads to regressions for most metrics. This algorithm also helps when applied to a non-instruction-tuned model, though it does not outperform supervised finetuning. “CSL” stands for “Constraint Satisfaction Level”. We show positive results (relative to the baseline) in blue, negative in orange, and neutral (within 0.5 for satisfaction rate or 0.05 for CSL) in gray; the top variant of a given model is bolded.
Our proposed approach, RLCF, demonstrates consistent gains across all benchmarks (Table 2, Table 3, and Table 4). On IFEval’s “loose” metrics (which apply minor preprocessing to responses before checking for correctness), RLCF improves Qwen-7B-Instruct by 2.8-3.0% (relative), as shown in the left half of Table 2. On FollowBench (shown in Table 3), RLCF achieves an 8.2% increase on Constraint Satisfaction Level (CSL; the expected proportion of constraints satisfied) and a 5.5% increase on average Hard Satisfaction Rate (how often all constraints are satisfied). RLCF also performs competitively on InFoBench (right half of Table 2), achieving results comparable to the best-performing reward model-based approaches while maintaining consistent gains across all constrain-based benchmarks. On “general use-case” instruction-following benchmarks, RLCF consistently increases the win rate of Qwen2.5-7B over GPT-4 (shown in Table 4), with the relative improvement ranging from 2.8% to 8.4%.
@@
Table 4: We compare methods on two “general” instruction-following benchmarks: Arena-Hard and AlpacaEval. RLCF gives modest but consistent gains on both the original metric and length/style-controlled metric on each benchmark. We show positive results (relative to the baseline) in blue, negative in orange, and neutral (within 0.5) in gray; the top variant of a given model is bolded.

### v2

@@ Abstract
Language models must be adapted to understand and follow user instructions. Reinforcement learning is widely used to facilitate this – typically using fixed criteria such as “helpfulness” and “harmfulness”. In our work, we instead propose using flexible, instruction-specific criteria as a means of broadening the impact that reinforcement learning can have in eliciting instruction following. We propose “Reinforcement Learning from Checklist Feedback” (RLCF). From instructions, we extract checklists and evaluate how well responses satisfy each item—using both AI judges and specialized verifier programs—then combine these scores to compute rewards for RL. We compare RLCF with other alignment methods on top of a strong instruction following model (Qwen2.5-7B-Instruct) on five widely-studied benchmarks – RLCF is the only method to help on every benchmark, including a 4-point boost in hard satisfaction rate on FollowBench, a 6-point increase on InFoBench, and a 3-point rise in win rate on Arena-Hard. We show that RLCF can also be used off-policy to improve Llama 3.1 8B Instruct and OLMo 2 7B Instruct. These results establish rubrics as a key tool for improving language models’ support of queries that express a multitude of needs. We release our our dataset of rubrics (WildChecklists), models, and code to the public. [^1] Code: www.github.com/viswavi/RLCF, Dataset: www.huggingface.co/datasets/viswavi/rlcf
@@ 1 Introduction
Language models must follow user instructions to be useful. As the general public integrates language model-based assistants into their completion of daily tasks, there is an expectation that models can faithfully follow the users’ requests, which increasingly involve rich, multi-step instructions (Liu2024WeNS; Zhao2024WildChat1C; zhenglmsys). Today’s models are almost universally trained to follow instructions via a two-step process: instruction finetuning, followed by reinforcement learning from human feedback (RLHF). Instruction finetuning, where the model is trained to mimic responses generated by annotators (Raffel2019ExploringTL), has historically been the primary workhorse for imbuing language models with some amount of instruction following ability (Wang2022SelfInstructAL; Chung2022ScalingIL; Xu2024MagpieAD; Lambert2024TLU3P). Model developers then frequently employ RLHF, where the model is trained to generate responses that look more like labeled “good” responses than “bad” responses, as a refinement step to decrease the likelihood that the model exhibits predefined poor behaviors (typically harmful behaviors) (Ziegler2019FineTuningLM; Bai2022ConstitutionalAH). Unlike “verifiable” tasks where reinforcement learning is a workhorse (DeepSeekAI2025DeepSeekR1IR; Lambert2024TLU3P; ifbench), reinforcement learning remains difficult to utilize for ambiguous or “non-verifiable” tasks, such as instruction following. What would it take to make RL a general-purpose solution at eliciting desirable behaviors in subjective or open-ended settings?
@@
Figure 1: RL on Checklist Feedback consistently improves Qwen2.5-7B-Instruct, whereas every other source of automatic feedback gives mixed results.
@@
We believe the solution must involve producing better reward signals. Recent work on RL for language model alignment has focused on automatically obtaining feedback on model behavior, either by (1) exclusively using instructions with verifiable answers (autoif; ifbench), (2) grading responses with specially-trained reward models (Wang2024HelpSteer2PreferenceCR; Eisenstein2023HelpingOH), or (3) distilling preferences from a larger prompted model (Bai2022ConstitutionalAH; Tunstall2023ZephyrDD). Using instructions with verifiable answers can only reinforce limited aspects of behavior (ignoring subjective constructs, e.g. topicality or style). Reward models (“RMs”) are flexible, but their notion of rewards can be arbitrary, leading to reward hacking (Eisenstein2023HelpingOH). When distilling preferences from a larger model, that model must infer what aspects to grade on, reducing the “generator-verifier gap” that enables RL (Swamy2025AllRL). Even if multiple criterion-specific prompts are used, these criteria may not be comprehensive (Bai2022ConstitutionalAH; Glaese2022ImprovingAO).
@@
In this paper, we ask: “how can we grade responses to instructions in a manner that is automatic (requires no human annotation), flexible (considers all aspects of response quality), intuitive (aligned with perceptible differences in responses), and applicable to any instruction or response, to enable more better language model alignment?” We propose extracting dynamic rubrics from instructions – an approach we term Reinforcement Learning from Checklist Feedback (RLCF). This approach reduces the task of grading responses to answering a series of yes/no questions, which can be answered by a model or by executing a verification program.
@@
On 5 benchmarks covering both constrained instruction following (IFEval, InFoBench, FollowBench) and general conversational assistance (AlpacaEval, Arena-Hard), we find that RLCF provides benefits on all instruction following benchmarks while maintaining improved performance on general conversational assistance benchmarks. In contrast, all alternative forms of AI feedback lead to mixed results, as shown in Figure 1. RLCF provides a 5.4% relative improvement over Qwen2.5-7B-Instruct in average hard satisfaction rate on FollowBench, a 6.9% relative improvement in overall requirement following ratio on InFoBench, and a 6.4% relative improvement on Arena-Hard (Jiang2023FollowBenchAM; qin2024infobench; ArenaHard). RLCF can also be used off-policy; we see Llama 3.1 8B Instruct and OLMo 2 7B Instruct improve using samples collected using Qwen2.5-7B-Instruct. Despite these considerable improvements, RLCF simply requires a teacher model, with no need for additional data or human annotations, making this approach amenable to diverse languages or domains. We provide evidence that checklist-based rewards are well-correlated to human preference judgments (comparable to many finetuned reward models) while providing a stronger learning signal than alternatives.
@@ 2 Checklist Generation
To satisfy definition #3, checklists must be comprehensive (cover most relevant aspects of quality) and natural (entailed by their corresponding instructions). Based on the observation that false positive rewards are often more detrimental to reinforcement learning than false negatives (Huang2024TheDS), we want checklists that are objective (facilitate automatic verification) and atomic (each requirement focuses on a single aspect of quality), to make requirement checking easier.
@@
  • Direct: We simply prompt an LM to extract a checklist from a given instruction (Cook2024TICKingAT). This approach is intuitive and simple but risks repeating the original instruction via these individual criteria, which may limit comprehensiveness and objectiveness.
@@
To compare these, we generate checklists for all instructions in InFoBench (qin2024infobench). We use gpt-4o to blindly evaluate each of these checklists on naturalness, objectivity, comprehensiveness, and atomicity, then select the better one overall. We manually perform the same evaluation on a subset of 50 instructions from the “Easy Set” of InFoBench.
@@
Regularization via universal criteria. In initial experiments, we found that optimizing for checklist completion led to responses beginning with long preamble overviews, suggesting reward hacking. Following Sun2023SALMONSW (who report a similar issue in prior work), we add two “universal requirements” to all generated checklists. These requirements state “1) The response directly address the request without excessive or off-topic information not necessary for addressing the user’s instruction? and 2) The response should match the context and the instruction, whether it requires professionalism, friendliness, formality, or neutrality.”, with a corresponding total importance weight of 100/100.
@@
Dataset Generation Using the candidate-based method, we generate checklists for 130,000 instructions from WildChat to create a new dataset, WildChecklists. To generate candidate responses for our method, we use Qwen2.5-0.5B, Qwen2.5-1.5B, Qwen2.5-3B, and Qwen2.5-7B (Yang2024Qwen25TR). Qwen2.5-72B-Instruct is the checklist generator model for both methods.
@@ 3 Reinforcement Learning from Checklist Feedback
Sampling Candidate Responses. To facilitate RL, we sample response pairs from our base policy with a temperature of 1.3 and a top-p of 0.9 (Holtzman2019TheCC). This is simpler than prior works that systematically perturb samples to induce greater complexity (conifer; autoif).
Flexible Scoring Given a prompt, a response, and an individual checklist item, we use a combination of an LM judge and a verifier program to grade the response. For each checklist item, we prompt a judge model (Qwen2.5-72B-Instruct) using the prompt in Appendix B to generate a numerical score between 0 and 100. We take the average of 25 numerical scores sampled from the model. [^2] We sample responses using the n parameter in vLLM (Kwon2023EfficientMM). This approach follows prior work that describes the importance of using the mean score rather than mode score from an LM-as-a-judge model (Wang2025ImprovingLI) Regardless, this makes the AI judge component the computational bottleneck of our pipeline. In Section 5.7, we show that n can be significantly reduced, at a modest accuracy cost. .
LLMs struggle to evaluate hard, discrete criteria, such as “does the response contain the letter R at least three times?” (Fu2024WhyDL). To handle such constraints, we follow prior work in generating a program to grade responses when needed (autoif; Zhou2023InstructionFollowingEF). We prompt a model (using the prompt in Appendix B) to write code only when the model is confident it can exactly check the requirement. For example, our program generator abstains from writing a verifier for the criterion “Is the sentence coherent”. The program-graded score is then averaged with the AI judge’s score. [^3] This approach is much simpler than the most relevant prior work that uses programs to evaluate responses, AutoIF (autoif), which uses test-case generation and LM-based filters to remove low-quality programs.
Preference Tuning. Given a separate numerical score for each criterion for each response, we take the average of these scores, weighted by the importance score of each criterion. We keep only the 40% of response pairs with the greatest difference along at least one criterion of its corresponding checklist. This removes response pairs that are too similar to offer useful pairwise signal. We then assign these responses as a preference pair for direct preference optimization (Rafailov2023DirectPO).
@@ 4.1 Experimental Details
Training Data As a fixed source of instructions for all methods, we use WildChat, a set of natural conversations between users and AI language models crowdsourced from users across the world (Zhao2024WildChat1C). We filter out conversations that are non-English, toxic, or longer than two turns.
@@
Training We finetune the model for 2 epochs using DPO with a batch size of 1024 and a maximum sequence length of 2048. We use a cosine learning rate schedule with a max LR of 3e-6 and a min LR of 2e-6.444When training models with Ultrafeedback, we instead used a minimum learning rate of 3e-7. We found this parameter resulted in a slightly stronger baseline when learning from this feedback. We use OpenRLHF for training (Hu2024OpenRLHFAE), and we train on one 8xH100 node with 80GB GPU memory, which took roughly 3 hours for each model.
Benchmark Data We evaluate our method on five benchmarks: IFEval (Zhou2023InstructionFollowingEF), InFoBench (qin2024infobench), FollowBench (Jiang2023FollowBenchAM), AlpacaEval (AlpacaEval), and Arena-Hard (ArenaHard). The first three of these measure instruction following ability in the presence of fine-grained constraints. The last two measure “general-purpose” instruction following ability, using naturalistic instructions based on user queries collected in the wild.
@@ 4.2 Baselines
Instruction Finetuning: We compare with instruction finetuning, to isolate the benefit of additional knowledge from the manner it is given (ground truth or rewards). Here, we distill (Hinton2015DistillingTK) from a larger model, Qwen2.5-72B-Instruct, finetuned via LlamaFactory (llamafactory).
Reward Models: We mirror our training approach for learning from checklist feedback, but using state-of-the-art reward models to decide which response should be chosen or rejected. We use Skywork/Skywork-Reward-Gemma-2-27B (Liu2024SkyworkRewardBO) and ArmoRM-Llama3-8B-v0.1 (Wang2024InterpretablePV) – both are highly rated on RewardBench (Lambert2024RewardBenchER). [^5] Skywork/Skywork-Reward-Gemma-2-27B and ArmoRM-Llama3-8B-v0.1 are ranked as #4 and #24, respectively, on RewardBench as of July 2025.
Prompted AI Judge: We compare against using the same “teacher” model as a judge, without using rubrics. We query this teacher in two settings: 1) “Ultrafeedback”, where the judge rates all candidate responses from 1-5 (Cui2023UltraFeedbackBL) separately across four quality aspects (instruction following, helpfulness, truthfulness, honesty) and averages these scores; and 2) AI Judge, where a near-identical prompt as RLCF is used (§3) to similarly sample 25 scores between 0 and 100 from the judge.
@@ 5.1 RL from Checklist Feedback consistently improves language models
Table 2: For instruction following, RLCF leads to large gains with open-ended constraints (InFoBench) and slightly positive or neutral changes on format-based constraints (IFEval). Off-the-shelf reward models help on InFoBench but hurt on IFEval. We show positive results (relative to the baseline) in blue, negative in orange, and neutral (within 0.5) in gray; the top variant of a given model is bolded.
@@
Table 3: RLCF leads to improvements on FollowBench across all metrics when starting with an instruction-tuned model, while using an off-the-shelf reward model for preference labeling leads to regressions for most metrics. This algorithm also helps when applied to a non-instruction-tuned model, though it does not beat supervised finetuning. “CSL” stands for “Constraint Satisfaction Level”. We show positive results (relative to the baseline) in blue, negative in orange, and neutral (within 0.5 for satisfaction rate or 0.05 for CSL) in gray; the top variant of a given model is bolded.
Our proposed approach, RLCF, demonstrates consistent gains across all benchmarks (Table 2, Table 3, and Table 4). On IFEval’s “loose” metrics (which apply minor preprocessing to responses before checking for correctness), RLCF improves Qwen-7B-Instruct by 2.8-3.0% (relative), as shown in the left half of Table 2. On FollowBench (shown in Table 3), RLCF achieves an 8.2% increase on Constraint Satisfaction Level (CSL; the expected proportion of constraints satisfied) and a 5.5% increase on average Hard Satisfaction Rate (how often all constraints are satisfied). RLCF also performs competitively on InFoBench (right half of Table 2), achieving results comparable to the best-performing reward model-based approaches while maintaining consistent gains across all constrain-based benchmarks. On “general use-case” instruction following benchmarks, RLCF consistently increases the win rate of Qwen2.5-7B over GPT-4 (shown in Table 4), with the relative improvement ranging from 2.8% to 8.4%.
@@
Table 4: We compare methods on two “general” instruction following benchmarks: Arena-Hard and AlpacaEval. RLCF gives modest but consistent gains on both the original metric and length/style-controlled metric on each benchmark. We show positive results (relative to the baseline) in blue, negative in orange, and neutral (within 0.5) in gray; the top variant of a given model is bolded.